# Skeleton: A New Framework for Accelerating Language Models via Task Neuron Localized Prompt Tuning

### v2

@@
Abstract
Prompt tuning methods have shown comparable performance to general training methods as parameter-efficient fine-tuning (PEFT) methods in various natural language understanding tasks. However, existing prompt tuning methods still utilize the entire model architecture even when solving a specific task, which prevents them from accelerating inference speed during the application procedure. In this paper, we propose a novel prompt tuning framework called Skeleton to efficiently utilize a language model in terms of memory and time complexity for solving various tasks, retaining only task-relevant neurons by using an explainability method. From our framework, we can efficiently solve various tasks by using only task-relevant neurons and prepending adequate task-specific prompt tokens with only a single language model. Experiments reveal that our method significantly enhances inference efficiency (at most ×$\times 1.73$ speed up) for various widely used benchmarks, showing comparable performance to the prompt tuning method. Moreover, our method is applicable across various transformer-based architectures, confirming its practicality and scalability.
Skeleton: A New Framework for Accelerating Language Models via Task Neuron Localized Prompt Tuning
Nakyeong Yang1, Jiwon Moon1, Junseok Kim2, Yunah Jang1 and Kyomin Jung1 1Seoul National University, 2Pohang University of Science and Technology {yny0506, wldnjs913, vn2209, kjung}@snu.ac.kr {junseokkim00}@postech.ac.kr
1 Introduction
With the success of parameter-efficient fine-tuning (PEFT) methods, pre-trained language models perform various tasks by adjusting only small numbers of parameters. Among PEFT methods, prompt tuning methods (Lester et al., 2021; Li and Liang, 2021; Liu et al., 2021) are widely used to solve multiple tasks since they require only a few tunable tokens to be prepended to the input tokens while freezing the original parameters of a language model (Chan et al., 2023; Huang et al., 2024). This strategy is memory efficient, as it enables a model to solve various tasks with only a few task-specific prompt tokens while retaining the original parameters. However, we can raise one research question regarding this task-specific trait of prompt tuning:
"Leveraging prompt tuning, could we condense task knowledge into a few task neurons in a frozen model to efficiently utilize it?"
@@ 1 Introduction
Figure 1: Skeleton Framework Overview. Our framework activates only task-localized paths of language models in the prompt tuning procedure. Each color (i.g., red, blue, and yellow) corresponds to a specific task; we obtain a task-specific prompt and neuron indices after executing our framework. Gray-colored paths in the figure represent task-irrelevant paths that are never used during training and inference, thereby improving inference speed.
Existing prompt tuning frameworks are limited by the need to use the full original model architecture during inference. In fact, these methods may require additional inference time, as the prepended prompt tokens introduce further computational overhead. Therefore, existing methods (Ma et al., 2022; Liang et al., 2023) have attempted to accelerate the inference speed of prompt-tuned models by pruning unnecessary prompt tokens. However, they have only marginally improved inference efficiency, as the number of pruned tokens is a small fraction ($\leq 0.1\%$) of all the parameters of a language model (Lester et al., 2021).
Prior studies have attempted to investigate and analyze task-relevant neurons in language models to explain the language models’ behavior (Panigrahi et al., 2023; Yang et al., 2024). Specifically, Panigrahi et al. (2023) has suggested training new parameters, which mask the original parameters of language models to determine task neurons. Yang et al. (2024) has utilized an explainability method to determine task-relevant neurons and reveal that some task-relevant neurons significantly affect the outputs of specific tasks (i.e., natural language understanding, social bias). However, while they have introduced a method to identify task-relevant neurons, they fall short of proposing a practical approach to efficiently utilize those neurons during training and inference.
To address this challenge, we propose a novel training framework, Skeleton, designed to enhance the efficiency of language models by leveraging only task-relevant neurons for solving various tasks. Specifically, we use attribution (Yang et al., 2023, 2024), an explainability method, to identify and utilize only task-relevant neurons in the prompt tuning procedure. Our framework requires only a few prompt tokens to be updated and freezes the original parameters of a language model, as it adopts the prompt tuning method. Furthermore, Skeleton dramatically accelerates the inference speed of language models since it utilizes only task-relevant neurons. For example, in order to solve a specific task (e.g., sentiment analysis, natural language inference), our method prepends prompt tokens related to the task and activates only the task-relevant neurons to solve the task.
@@
We demonstrate our method in various widely used natural language understanding benchmarks and reveal that our method significantly outperforms existing PEFT methods in inference efficiency without sacrificing the original performance of each task. Surprisingly, our method improves the inference speed up to ×$\times 1.73$ with 65% of the parameters pruning rate. We additionally conduct various analyses to offer a guideline for using our new prompt tuning framework, such as investigating which types of layers are crucial for processing task knowledge in a language model. Furthermore, our prompt tuning framework is practical and scalable since it is applicable to any transformer-based architecture. In summary, this work makes the following contributions:
  • We first define task neuron and propose Task Neuron Localized Prompt Tuning (Skeleton), a practical training framework for accelerating the inference speed of a language model for various tasks.
@@
  • We reveal that Skeleton successfully enhances the inference speed of each task without compromising the task performance.
2 Backgrounds
2.1 Parameter-efficient Fine-tuning
@@
Parameter-efficient Fine-tuning (PEFT) methods have emerged as a standard of efficient training methods for language models (Lester et al., 2021; Li and Liang, 2021; Liu et al., 2021; Hu et al., 2021; Poth et al., 2023). Among them, prompt tuning methods (Lester et al., 2021; Li and Liang, 2021; Liu et al., 2021) have been adopted for various NLP tasks (Wang et al., 2023; Bansal et al., 2023). Prompt tuning methods only update prompt tokens while maintaining the original parameters of a language model; thus, this characteristic makes the training process memory-efficient and effective in injecting task knowledge into a language model. However, they still face a limitation as they are not parameter-efficient during the inference process. Therefore, existing methods (Ma et al., 2022; Liang et al., 2023) propose prompt token selection methods to enhance the inference speed of prompt-tuned language models. Nevertheless, they have only marginally enhanced inference efficiency since the number of pruned tokens is only a tiny proportion of the entire number of input tokens.
2.2 Task Knowledge Localization
In spite of the significant performance of language models, it remains challenging to accurately investigate the role of each parameter in language models in dealing with a specific task. Therefore, existing studies have attempted to identify task-specific neurons in executing a specific task (Panigrahi et al., 2023; Yang et al., 2023, 2024). Panigrahi et al. (2023); Yang et al. (2023, 2024) have studied to localize neurons for performing a specific task by clarifying the role of each parameter in a language model. Panigrahi et al. (2023) has suggested a training-based task knowledge localization method, model grafting, which localizes a task knowledge by training new parameters to mask original parameters. Although it has effectively identified task neurons, it has required excessive additional storage for newly trained masking parameters, which is equivalent to the number of model parameters. Yang et al. (2023, 2024) has illuminated task neurons by utilizing the attribution technique (Shrikumar et al., 2016), an explainability method that derives the importance of each feature when solving a specific task. Yang et al. (2023, 2024) has verified that the attribution effectively detects task neurons for solving a specific task and proposed a task neuron detection method applicable to language modeling tasks. However, Yang et al. (2023, 2024) has a limitation in just pruning the task neurons; if we prune a large number of neurons, the performance of a specific task can be significantly degraded. Furthermore, the prior study has only proposed the method for pruning task neurons and is far from proposing a practical method to improve the inference speed of task-localized language models. The architecture of transformer variants consists of complex neural computation processes; thus, satisfying the integrity of dimension dependency for adjacent networks while pruning the weights is challenging. We adopt the attribution method (Yang et al., 2023, 2024) and propose a new practical framework for accelerating a language model, where the framework utilizes only task neurons while training a language model via prompt tuning, resolving the challenges in pruning networks in transformer variants.
3 Problem Definition
3.1 Prompt Tuning
@@
We adopt the prompt tuning method to propose a new framework for accelerating a language model. Formally, suppose we have a function $\mathcal{P}_{\theta}(y|x)=\prod_{k=1}^{K}\mathcal{P}_{\theta}(y_{k}|x,y_{1},...,y_{k-1})$ that represents a language model, where $\theta$ is parameters of the language model. Prompting is the method of adding extra information to the model to condition during its generation of $y$. Normally, prompting is achieved by adding a series of tokens $P_{\tau}\in\mathbb{R}^{l\times d}$ to the input $x$ as $\mathcal{P}_{\theta}(y|[P_{\tau};x])$, where $l$ and $d$ are token length and feature dimension, and $P_{\tau}$ is trainable parameters for a specific task $\tau$. In the prompt tuning process, we update only $P_{\tau}$ and get $P^{\prime}_{\tau}$ to maximize the likelihood of $y$ using gradient descent while freezing the parameters of the language model $\theta$. After the prompt tuning, we get the new conditional generation function $\mathcal{P}_{\theta}(y|[P^{\prime}_{\tau};x])$.
3.2 Skeleton Framework
In the Skeleton framework, we still update only prompt tokens, freezing the original parameters. However, we assume that there is a minimal parameter combination $\theta_{\tau}\subset\theta$ for solving a specific task $\tau$; thus, we first define task neuron as follows:
Task Neuron.
@@ 3.2 Skeleton Framework
Let $\mathcal{P}_{\theta}$ be a language model, where $\theta=\{\theta_{1},...,\theta_{N}\}$ is a set of parameters of it. Then, $\theta_{\tau}\subset\theta$ is defined as task neurons for a task $\tau$ if the below equation is satisfied:
$\sum_{(x,y)\in\mathcal{D}_{\tau}}\mathcal{L}(y,\mathcal{P}_{\theta}([P^{\prime}_{\tau};x]))\approx\sum_{(x,y)\in\mathcal{D}_{\tau}}\mathcal{L}(y,\mathcal{P}_{\theta_{\tau}}([P^{\prime\prime}_{\tau};x]))$ (1)
@@
where $\mathcal{D}_{\tau}$ means a dataset for a task $\tau$ and $\mathcal{L}$ is a score function (e.g., loss function or accuracy). $P^{\prime}_{\tau}$ and $P^{\prime\prime}_{\tau}$ are task-specific prompts. $N$ means the total number of parameters in a language model.
@@
Skeleton framework adopts the prompt tuning strategy using only task neurons $\theta_{\tau}$, and get the new task-localized conditional generation function $\mathcal{P}_{\theta_{\tau}}(y|[P^{\prime\prime}_{\tau};x])$. Therefore, we can utilize the language model more time-efficiently by using only those neurons within the Skeleton framework. In addition to being time efficient, this framework is also memory-efficient since it requires only a small fraction of original parameters $\theta_{\tau}$ and just a few task-specific prompt tokens $P^{\prime\prime}_{\tau}$.
4 Methods
@@
In this section, we describe the process of Skeleton. We first train a language model using the prompt tuning method to inject prior knowledge of a specific task into it. After that, we quantify the task relevance of each neuron to determine task neurons in a language model. We finally retrain the prompt tokens with the prompt tuning method to condense the task knowledge to only the task neurons. Note that the original language model maintains only task neurons via a pruning method in this process. After applying Skeleton framework, we can utilize the language model efficiently by using only the task-specific prompt and task neurons, accelerating the inference speed for a specific task. The entire process of Skeleton is described in Figure 2 and Algorithm 1.
@@
Figure 2: Task Neuron Localization Process of Skeleton. After training a model using our framework, a task prompt and neurons are assigned to each task, and we utilize only those task-specific resources in the inference procedure. Red-colored values in the prune weights, $\tilde{W}_{1}$ and $\tilde{W}_{2}$, mean task neurons for a specific task (task A). The weight of FFN2 is pruned depending on FFN1’s dimension, and the output representation of FFN2 is zero-recovered (Blue-colored values) to match the output dimension of the residual network (described in Section 4.5).
4.1 Quantifying Task Relevance of Neurons

### v3

# Unplug and Play Language Models: Decomposing Experts in Language Models at Inference Time

@@
Abstract.
Enabled by large-scale text corpora with huge parameters, pre-trained language models operate as multi-task experts using a single model architecture. However, recent studies have revealed that certain neurons play disproportionately important roles in solving specific tasks, suggesting that task-relevant substructures can be isolated and selectively activated for each task. Therefore, we introduce Decomposition of Experts (DoE), a novel framework that dynamically identifies and activates task-specific experts within a language model to reduce inference cost without sacrificing accuracy. We first define a task expert as a set of parameters that significantly influence the performance of a specific task and propose a four-step unplug-and-play process: (1) receiving a user request, (2) identifying the corresponding task expert, (3) performing inference using the expert-localized model, and (4) restoring the original model and waiting for the next task. Using attribution methods and prompt tuning, DoE isolates task-relevant neurons, minimizing computational overhead while maintaining task performance. We assume a setting where a language model receives user requests from five widely used natural language understanding benchmarks, processing one task at a time. In this setup, we demonstrate that DoE achieves up to a $\times 1.73$ inference speed-up with a 65% pruning rate, without compromising accuracy. Comparisons with various task expert localization methods reveal that DoE effectively identifies task experts, while ablation studies validate the importance of its components. Additionally, we analyze the effects of batch size, token count, and layer types on inference speed-up, providing practical insights for adopting DoE. The proposed framework is both practical and scalable, applicable to any transformer-based architecture, offering a robust solution for efficient task-specific inference.
@@ Abstract.
Figure 1. Decompostion of Experts (DoE) Framework Overview. Our framework leverages an unplug-and-play strategy by identifying and activating only the task expert of a language model during the inference procedure when given a user request. Each color (i.e., red, blue, and yellow) corresponds to a specific task; we utilize a task-specific prompt and neuron indices obtained through a prompt tuning process. Gray-colored paths in the figure represent task-irrelevant paths that are never used during training and inference, thereby improving inference speed.
Knowledge Localization, Decomposition of Experts, Language Model
journalyear: 2025copyright: noneconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Korea.booktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreaisbn: 979-8-4007-2040-6/2025/11doi: 10.1145/3746252.3761011ccs: Computing methodologies Natural language processing
1. Introduction
Training on large-scale text corpora with numerous parameters enables recent pre-trained language models to perform a wide range of tasks using a single model architecture (Devlin et al., 2018; Liu et al., 2019; Lewis et al., 2019; Touvron et al., 2023; OpenAI, 2023). In this context, previous studies have supposed that each neuron in a language model may have a distinct role; thus, they have tried to investigate the role of individual neurons to explain task-specific knowledge in language models (Panigrahi et al., 2023; Yang et al., 2024; Chen et al., 2025). They have addressed various tasks, including sentiment analysis, natural language inference, and social bias detection. Therefore, this line of studies has laid the groundwork for identifying task-relevant neurons in language models by validating that these neurons possess sufficient knowledge to solve a specific task.
Based on these prior studies, we propose a novel framework, Decomposition of Experts (DoE), that improves the efficiency of language models by dynamically selecting task-specific experts in response to user requests. This approach reduces inference time in single-task scenarios. We first define a task expert as a subset of parameters that significantly contribute to solving a specific task. Based on this definition, we describe the DoE framework as an unplug-and-play language model composed of four stages: (1) receiving a user request, (2) identifying the corresponding task expert, (3) performing inference using the expert-localized model, and (4) restoring the original model and waiting for the next task (see Figure 1). Specifically, we employ the attribution method (Yang et al., 2024) to identify and activate only task-relevant neurons, which we treat as the task expert. DoE also adopts prompt tuning (Lester et al., 2021) to align task knowledge more effectively within parameters, enabling expert selection using a minimal number of neurons.
While previous pruning approaches have primarily focused on permanently removing unnecessary neurons or connections to reduce the overall parameter count, they typically rely on fixed subnetwork architectures optimized for specific tasks or consistently utilize lightweight models during inference (Frankle and Carbin, 2018; Ma et al., 2023; Chen et al., 2024). In contrast, our method assumes that pre-trained language models inherently possess sufficient multi-task capability. Rather than statically removing parameters, we achieve efficiency by dynamically selecting task-specific expert neurons and temporarily deactivating irrelevant ones (unplug). These neurons can then be reactivated for different tasks, enabling a reversible and task-adaptive computation flow. From this perspective, our framework offers a novel efficiency strategy that lies at the intersection of pruning and modularization. It is particularly advantageous in multi-task settings where rapid task switching is essential. DoE dramatically accelerates the inference speed of language models since it uses only a task expert, which is a small part of the entire model during inference. For example, to solve a specific task (e.g., sentiment analysis, natural language inference), our method prepends prompt tokens related to the task and activates only a task expert to resolve the task.
@@ 1. Introduction
Our experimental settings assume that a language model is given user requests about five widely used natural language understanding benchmarks. The experimental results demonstrate that our method significantly enhances inference efficiency without sacrificing the original performance of each task. Surprisingly, our method improves the inference speed by up to $\times 1.73$ with a 65% parameter pruning rate while solving a specific task. In addition, we conduct experiments to compare various task expert selection methods and reveal that the attribution method effectively identifies task experts. We also investigate the effects of batch size, the number of tokens, and types of layers on inference speed-up. Our framework is practical and scalable, as it can be applied to any transformer-based architecture. In summary, this work makes the following contributions:
  • We first define task expert and propose Decomposition of Expert (DoE), a novel language model usage framework that accelerates inference speed by identifying and isolating a task expert within a language model, leveraging an unplug-and-play strategy.
@@
  • We demonstrate that DoE effectively identifies task experts over other task expert localization methods and enhances the inference speed of each task without compromising task performance.
  • We offer various practical guidelines for adopting DoE by investigating the effects of hyper-parameters on inference speed-up and evaluating the scalability of the method.
2. Backgrounds
2.1. Task Knowledge Localization
Prior works have shown that language models trained with a vast size of corpora with a huge number of parameters can serve as a multi-task expert (Radford et al., 2019; Raffel et al., 2020). These multi-tasking abilities are attributed to exploiting commonalities and differences recognized while training multiple categories of texts (Zhang et al., 2022; Chen et al., 2024). Recent studies have shown that multi-task knowledge is distributed across sparse neurons in language models (Liu et al., 2023; Chan et al., 2023). Building on this observation, earlier research has investigated the roles of individual parameters in handling specific tasks (Panigrahi et al., 2023; Chen et al., 2025; Yang et al., 2024).
Panigrahi et al. (2023) have suggested a training-based task knowledge localization method, model grafting, which identifies task-relevant neurons by training new parameters to mask original parameters. Chen et al. (2025) have investigated the query-relevant neurons by utilizing intergrated gradients (Sundararajan et al., 2017), an explainability method that derives the importance of each neuron when inferring a long text. Yang et al. (2024) have illuminated bias-relevant neurons by utilizing attribution (Shrikumar et al., 2016), an explainability method that derives the importance of each neuron when inferring a specific biased text. Yang et al. (2024) have verified that the attribution effectively detects bias-relevant neurons and proposed a method applicable to language modeling.
Although Panigrahi et al. (2023); Chen et al. (2025); Yang et al. (2024) have effectively identified task-relevant neurons, they have not proposed practical methods to enhance the usability of language models. Panigrahi et al. (2023) requires excessive additional storage and computation for newly trained masking parameters, equivalent to the number of model parameters. In addition, Yang et al. (2024) has a limitation in just eliminating the knowledge-relevant neurons; if we prune a large number of neurons, the performance of a specific task can be significantly degraded. In other words, this line of studies has only focused on investigating knowledge-relevant neurons, failing to propose a practical method for efficiently utilizing language models during the inference procedure.
2.2. Language Model Efficiency
Existing studies have proposed various approaches to enhance the efficiency of language models, which can be categorized into three main groups: (1) knowledge distillation, (2) quantization, and (3) pruning. Knowledge distillation transfers the knowledge of larger or stronger models into smaller or more efficient ones (Gu et al., 2023; Agarwal et al., 2024). Quantization reduces the precision of model parameters and activations to lower memory and computational costs (Shao et al., 2023; Yue et al., 2024; Xu et al., 2024). Pruning eliminates less important parameters to reduce model size and inference time (Frankle and Carbin, 2018; Ma et al., 2023; Chen et al., 2024).
While these methods have improved efficiency, they typically rely on statically modifying the model into fixed architectures tailored for specific tasks and deploying task-specific lightweight models consistently at inference time. In contrast, our approach proposes a new framework of language model usage: Unplug-and-plug language models: preserving the full model and dynamically activating only task-relevant neurons, enabling adaptive and reversible task-specific inference without architectural reconfiguration.
2.3. Parameter-efficient Fine-tuning.
@@ 2.2. Language Model Efficiency
Parameter-efficient Fine-tuning (PEFT) methods have emerged as a standard of efficient training methods for language models (Lester et al., 2021; Li and Liang, 2021; Liu et al., 2021; Hu et al., 2021; Poth et al., 2023). Among them, prompt tuning methods (Lester et al., 2021; Li and Liang, 2021; Liu et al., 2021) have been adopted for various natural language processing tasks (Wang et al., 2023; Bansal et al., 2023). Prompt tuning methods only update prompt tokens while maintaining the original parameters of a language model; thus, this characteristic makes the training process memory-efficient and effective in injecting task knowledge into a language model. However, they are still limited as they are not parameter-efficient during the inference process. Therefore, existing methods (Ma et al., 2022; Liang et al., 2023) have attempted to accelerate the inference speed of prompt-tuned models by pruning unnecessary prompt tokens. However, they have only marginally improved inference efficiency, as the number of pruned tokens is a small fraction ($\leq 0.1\%$) of all the parameters of a language model (Lester et al., 2021). Our DoE framework adopts the prompt tuning method (Lester et al., 2021) to align the original task knowledge when localizing a task expert.
3. Problem Definition
3.1. Unplug-and-play language model
We define unplug-and-play language model as a language model that satisfies the following two conditions: (1) it should perform a single task using only a tiny fraction of its parameters, and (2) the model architecture is fully reversible; thus, it can restore its original parameters after task localization, allowing the model to remain idle to be available for other tasks. Typical compression methods, such as pruning and knowledge distillation, cannot be considered unplug-and-play language models since they statically modify the model into fixed architectures for specific tasks and consistently deploy task-specific lightweight models during inference (irreversible).
3.2. Prompt Tuning
@@
We adopt the prompt tuning method to propose a new framework for efficiently utilizing a language model. Formally, suppose we have a function $\mathcal{P}_{\theta}(y|x)$ that represents a language model, where $\theta$ is the parameters of the language model. Prompting is the method of adding extra information to the model to condition during its generation of $y$. Normally, prompting is achieved by adding a series of tokens $P_{\tau}\in\mathbb{R}^{l\times d}$ to the input $x$ as $\mathcal{P}_{\theta}(y|[P_{\tau};x])$, where $l$ and $d$ are token length and feature dimension, and $P_{\tau}$ is trainable parameters for a specific task $\tau$. In the prompt tuning process, we update only $P_{\tau}$ and get $P^{\prime}_{\tau}$ to maximize the likelihood of $y$ using gradient descent while freezing the parameters of the language model $\theta$. After the prompt tuning, we get the new conditional generation function $\mathcal{P}_{\theta}(y|[P^{\prime}_{\tau};x])$. However, we assume that there is a minimal parameter combination $\theta_{\tau}\subset\theta$ for solving a specific task $\tau$; thus, we define task expert as follows:
3.3. Task Expert
@@
Let $\mathcal{P}_{\theta}$ be a language model, where $\theta=\{\theta_{1},...,\theta_{N}\}$ is a set of parameters of it. Then, $\theta_{\tau}\subset\theta$ is called as a task expert for a task $\tau$ if the below formula is satisfied:
(1) $\scriptsize\sum_{(x,y)\in\mathcal{D}_{\tau}}\hskip-9.10509pt\mathcal{L}(\mathcal{P}_{\theta},x,y)=\epsilon+\hskip-8.5359pt\sum_{(x,y)\in\mathcal{D}_{\tau}}\hskip-9.10509pt\mathcal{L}(\mathcal{P}_{\theta_{\tau}},x,y)\small$
@@
where $\mathcal{D}_{\tau}$ denotes the dataset for a task $\tau$ and $\mathcal{L}$ is a score function (e.g., loss function or accuracy) evaluated on the given model. $N$ means the total number of parameters in the language model. $\epsilon$ is a small error value. In other words, if the scores computed on $\mathcal{P}_{\theta}$ and $\mathcal{P}_{\theta_{\tau}}$ for the given dataset $\mathcal{D}_{\tau}$ are approximately the same, we can utilize only the task neurons $\theta_{\tau}$ to solve the task $\tau$.
@@
DoE framework adopts the prompt tuning strategy using only the task expert $\theta_{\tau}$ in a language model; thus, we get the task-localized conditional generation function $\mathcal{P}_{\theta_{\tau}}(y|[P^{\prime}_{\tau};x])$ instead of $\mathcal{P}_{\theta_{\tau}}(y|x)$. Notice that since the prompt tuning strategy does not modify the original parameters of language models, a model that adopts the prompt tuning can still be used as a unplug-and-play language model. We can utilize the language model more time-efficiently by using only the sparse neurons within our  framework. In addition to being time-efficient, this framework is also memory-efficient since it requires only a small fraction of the original parameters $\theta_{\tau}$ and just a few task-specific prompt tokens $P^{\prime}_{\tau}$.
4. Decomposition of Expert (DoE)
@@
In this section, we describe the process for adopting the DoE framework. We first train a language model using the prompt tuning method to align the prior knowledge of a specific task. After that, we quantify the task relevance of each neuron to determine a task expert in a language model. Finally, we apply prompt tuning to retrain the prompt tokens, enabling the condensation of task knowledge into the task expert. During this process, only task-relevant neurons are preserved in the original language model. The method is performed for a set of predefined $T$ tasks; thus, we have $T$ task-specific prompts and $T$ sets of task experts. After applying the DoE framework, we can utilize the language model efficiently by using only the task-specific prompt and task expert, accelerating the inference speed for a specific task. The entire process of DoE is described in Figure 2 and Algorithm 1.
@@
Figure 2. Task Expert Localization Process of DoE. After training a model with the prompt tuning, a task prompt and neurons are assigned to each task, and we utilize only those task-specific resources in the inference procedure. Red-colored values in the prune weights, $\tilde{W}_{1}$ and $\tilde{W}_{2}$, mean task neurons for a specific task (task A). The weight of FFN2 is pruned depending on FFN1’s dimension, and the output representation of FFN2 is zero-recovered (Blue-colored values) to match the output dimension of the residual network (described in Section 4.5).
4.1. Task Relevance Quantification